OpenStudy (anonymous):

HELP AGAIN ! Leonard owns a business earning $37,800 in profits. These profits are appreciating at about 2.3% each year. What are Leonard’s profits after eight years?

OpenStudy (anonymous):

CORRECT ME !

OpenStudy (anonymous):

HOLD ON !

OpenStudy (anonymous):

6955.2

OpenStudy (anonymous):

44755.2 my bad its this one

OpenStudy (anonymous):

A=(PRT)+P A=(37800*0.023*8) +37,800 A=6955.2+37,800 A=44,775.2